how can I get the "real GDP per capital" and "real GDP per hour" ?

Year

Population, annual average of mid-quarter estimates (thousands) Nominal GDP at market prices, billions of current CAD$ CPI, 2007=100 Number of Jobs (thousands) Annual average hours per worker Working Age Population (thousands) Labour Force (thousands) Employment (thousands) GDP PPP exchange rate, US$/CAD$
1990 27632.4 693.0 70.3 13410.5 1,797 18334.5 14244.6 13086.4 0.81

Please tell me what to enter in EXCEL to calculate the answer I need. Pay attention to the conversion unit.
Answer #1

Solution:

For real GDP per capita-

Nominal GDP in us $= Nominal GDP in CAD $*0.81=693000 million*0.81= us  $561330 million.

Nominal GDP per capita= Nominal GDP/ Population=561330 million/27632400= us $20314.19

Real GDP Per capita =(​CPI in current year/ CPI in base year)*Nominal GDP per capita=(​70.3/10O)*20314.19= us $14280.87

for real GDP per hour-

no. of hours= number of Jobs* Average annual hours per worker=13410500*1797=2410675500 hours

nominal GDP per hour =_ Nominal GDP/ No. of hours=561330 million/2410675500= us $232.85

Real GDP per hour =(70.3/100)*232.85=_us $163.69

Above solution can be inserted into excel accordingly.